Find the thevinin resistance of the circuit. I got 118.46 ohms and would like to check my answer. Thanks

R1 R5 M Vout+ 15 100 V1 + R4 R3 470 5 220 R2 Vout- 220
Answer #1

Answer -

   For calculating Thevenin resistance, we have to short circuit the DC voltage source V1 = 5 V.

After that the total resistance as seen from the output terminal is ,

   \small R_{th}=\left [ \left [ [R_{1}+R_{2}]\left | \right |R_{3} \right ]+R_{5} \right ]\left | \right |R_{4}

   \small R_{th}=\left [ \left [ [15+220]\left | \right |470 \right ]+100 \right ]\left | \right |220

   \small R_{th}=\left [ \left [ 235\left | \right |470 \right ]+100 \right ]\left | \right |220

   \small R_{th}=\left [ 156.67+100 \right ]\left | \right |220

   \small R_{th}= 256.67\left | \right |220

     \small R_{th}= 118.46  \small \Omega

So, the answer is right.